# Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs

### v2

@@ 3.2.1 The Lovász theta relaxation
The Lovász theta relaxation has been introduced in 1979, yet we refer to the presentation of [43]. Introduce the quadratic variable $x_{ij}$, representing the product $x_{i}x_{j}$ for all $\{i,j\}\in V\times V$ and let $X=xx^{\top}$ be the associated symmetric matrix of order $n=|V|$. An upper bound to $\alpha(G)$ is given by
$\theta(G,w)=\max$ $\;\sum_{i\in V}w_{i}x_{i}$ (SDP-$\theta$)
s.t. $\;X_{ii}=x_{i},\quad i\in V,$
@@
$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$ (10)
We simply denote it by $\theta(G)$ in the unweighted case. The projection $\operatorname{TH}(G)$ of the feasible region of (SDP-$\theta$) onto the subspace of the $x_{i}$ variables is known as the theta body, a convex but not polyhedral set in general. Remarkably, Grötschel et al. [33] proved that $\operatorname{STAB}(G)\subseteq\operatorname{TH}(G)\subseteq{\operatorname{QSTAB}(G)}$, where equality holds if and only if $G$ is perfect. This implies that there exists a polynomial-time separation algorithm for a class of inequalities which includes all clique inequalities. A thorough comparison between the original formulation in [42] and (SDP-$\theta$) is illustrated in [26]. $\theta(G)$ can be computed in polynomial time up to an arbitrary precision and often provides a strong bound to the stability number $\alpha(G)$, typically better than those obtained from linear relaxations see, e.g., [38, 69, 28, 30]. Another interesting feature of (SDP-$\theta$) concerns with its computational behavior. In fact, (SDP-$\theta$) is handled better than similarly sized unstructured SDPs by general SDP solvers. Moreover, several algorithms have been designed to solve it efficiently as in, e.g., [54, 44, 29]. Therefore, $\theta(G)$ achieves a good compromise between strength of the upper bound and computational burden. From a theoretical viewpoint, Busygin and Pasechnik [13] showed that, unless P=NP, no polynomially computable upper bound to $\alpha(G)$ which is provably smaller than $\theta(G)$ can be found. In this context, several stronger relaxations have been obtained by adding linear inequalities to (SDP-$\theta$), as reviewed in the following subsections.
@@ 3.2.2 The Schrijver relaxation
Schrijver [59] observed that the PSD condition (10) does not imply the non-negativity on $X$. Hence, by adding the inequalities
$X_{ij}\geq 0,\quad\forall\{i,j\}\in\bar{E},$ (SDP-$\theta^{+}$)
to (SDP-$\theta$), one yields a model that we denote as (SDP-$\theta^{+}$) and the corresponding upper bound as $\theta^{+}(G)$. Accordingly, $\operatorname{TH}^{+}(G)$ denotes the convex set obtained by its projection onto the subspace of $x_{i}$ variables.
@@ 3.2.3 The Lovász-Schrijver relaxation
The Lovász-Schrijver relaxation for the SSP is obtained by applying the operator $N_{+}(\cdot)$ to $\operatorname{FRAC}(G)$. $N_{+}(\operatorname{FRAC}(G))$ corresponds to (SDP-$\theta^{+}$) plus the following linear inequalities:
$X_{ij}$ $\geq 0,$ $\{i,j\}\in\bar{E},$ (11)
$X_{ik}+X_{jk}$ $\leq x_{k},$$\{i,j\}\in E,k\neq i,j,$ (12)
$x_{i}+x_{j}+x_{k}$ $\leq 1+X_{ik}+X_{jk},$$\{i,j\}\in E,k\neq i,j.$ (13)
Lovász and Schrijver [43] assessed the theoretical strength of $N_{+}(\operatorname{FRAC}(G))$ by showing that it satisfies several well-known classes of valid inequalities for $\operatorname{STAB}(G)$. Namely, they proved that $N_{+}(\operatorname{FRAC}(G))$ satisfies all clique, odd cycle, odd antihole and odd wheel inequalities. Giandomenico and Letchford [27] showed that, in addition, it satisfies all web inequalities. Optimizing $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ returns an upper bound to $\alpha(G)$ that we denote as $\lambda(G)$. From a more general perspective, it is interesting to observe the key role of the positive semidefiniteness constraint. In fact, Burer and Vandenbussche [12] compared $\lambda(G)$ with the optimal value of the corresponding LP relaxation obtained by the Sherali-Adams procedure, observing a substantial improvement due to it.
Handling such a strong relaxation in practice is not straightforward. The first related computational results, obtained by a lift-and-project cutting plane method, appear in [3]. Results on optimising directly over $M_{+}(\operatorname{FRAC}(G))$ are illustrated in [21] and, more recently, in [12] using a specialised augmented Lagrangian method. The resulting upper bounds to $\alpha(G)$ may be considerably better than $\theta^{+}(G)$ but at the expense of a large increase in running times. In practice, this relaxation turns out to be hardly tractable for graphs of medium/large size.
@@ 3.2.4 Further SDP relaxations
Variants of the previous relaxations are presented in [34, 23]. In the former, extra triangle inequalities are added to $N_{+}(\operatorname{FRAC}(G))$ to obtain a very strong relaxation. In the latter, relaxations are specialised with respect to graph density to improve efficiency. We refer to [26] for an exhaustive review of these methods. Two SDP-related methodologies are also presented in the literature. In [24] strong SDP relaxations of several combinatorial optimization problems are investigated by exploiting the exact subgraph inequalities. Despite the power of this technique, the experience reveals that strengthening the natural SDP relaxation of the SSP is typically harder than other problems, e.g., max-cut, and confirms that improving the $\theta$ bound towards $\alpha(G)$, even by a small amount, is actually challenging. Another effective method (presented in the context of the clique number) based on adding non-valid inequalities has been developed in [41]. This turns out to achieve quite strong bounds for both structured and random graphs with up to 150 vertices.
@@ 4 New SDP relaxations
$x_{i}x_{j}\rightarrow$ $\quad X_{ij}\geq 0,$ (14)
$x_{i}(1-x_{j})\rightarrow$ $\quad x_{i}\geq X_{ij},$ (15)
$(1-x_{i})x_{j}\rightarrow$ $\quad x_{j}\geq X_{ij},$ (16)
$(1-x_{i})(1-x_{j})\rightarrow$ $\quad x_{i}+x_{j}-X_{ij}\leq 1.$ (17)
The linear inequalities (14)–(17) are known as McCormick [46] inequalities and describe the convex hull of the set $\{(x_{i},x_{j},y_{ij})\in\{0,1\}^{3}:y_{ij}=x_{i}x_{j}\}$. It is easy to see that they are not implied by the PSD condition (6) and should be explicitly added to the relaxations illustrated below unless implied by other constraints.
@@ 4.1 Relaxation $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$
In order to investigate $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ we first describe the application of the operator $M_{+}(\cdot)$ to $\operatorname{QSTAB}(G)$. The lifting step consists in multiplying each clique inequality $\sum_{i\in C}x_{i}\leq 1$, $C\in{\cal K}$, by $x_{j}$ and $(1-x_{j})$, for all $j\in V$. Then, adding the condition $X_{ii}=x_{i}$, one obtains:
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C\setminus\{j\}}X_{ij}\leq 0,$ if $j\in C$, (18)
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C}X_{ij}-x_{j}\leq 0,$ if $j\not\in C$, (19)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}x_{i}-\sum\limits_{i\in C\setminus\{j\}}X_{ij}\leq 1,$ if $j\in C$, (20)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}\leq 1,$ if $j\not\in C$. (21)
Inequalities (19), called clique-variable inequalities, have been introduced in [28]. One can observe that
@@
    Constraints (18) imply $X_{ij}=0$ for all $\{i,j\}\in E$ and constraints (20) reduce to the corresponding clique inequality $\sum_{i\in C}x_{i}\leq 1,C\in{\cal K}$, which is already implied from (19) and (21) by Remark 1;
@@
    $X_{ij}=0$ for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (19) imply inequalities (15) and (16);
@@
$\begin{split}{M_{+}(\operatorname{QSTAB}(G))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ \sum\limits_{i\in C}X_{ij}-x_{j}&\leq 0,\quad C\in\mathcal{K},j\in V\setminus C,\\ \sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}&\leq 1,\quad C\in\mathcal{K},j\in V\setminus C,\\ x_{i}&=X_{ii},\quad i\in V,\\ X_{ij}&=0,\quad\{i,j\}\in E,\\ X_{ij}&\geq 0,\quad\{i,j\}\in{\bar{E}}\\ x_{i}&\geq 0,\quad i\in V\bigg{\}}.\end{split}$ (22)
Moreover, one has:
@@
We denote by $\mu(G)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{QSTAB}(G))}\right\}$ the resulting upper bound to $\alpha(G)$. The following example shows that the containment ${M_{+}(\operatorname{QSTAB}(G))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ can be strict.
@@
The example indeed suggests a general result. In fact, we can show that, unlike $N_{+}(\operatorname{FRAC}(G))$, relaxation $N_{+}(\operatorname{QSTAB}(G))$ implies the class of antiweb inequalities.
@@
$N_{+}(\operatorname{QSTAB}(G))$ implies antiweb inequalities
Proof Let $G$ be the $(p,q)$-antiweb graph with $V=\left\{1,\ldots,p\right\}$. In what follows, we use arithmetic modulo $p$. Note that, by definition, any consecutive $q$ nodes in $G$ form a clique. By denoting $\alpha=\left\lfloor\frac{p}{q}\right\rfloor$, the $(p,q)$-antiweb inequality is
@@
Given any $i\in V$, we denote with $G\setminus i$ the subgraph induced by the vertex set
@@
is valid for ${\operatorname{QSTAB}(G\setminus i)}$, for all $i\in V$, then the theorem’s statement will follow from Lemma 1. Let $i$ be any node in $V$ and consider the following $\alpha-1$ cliques in $G\setminus i$:
$C^{1}$ =$=\left\{i+q,\ i+q+1,\dots,\ i+2q-1\right\},$
$C^{2}$ =$=\left\{i+2q,\ i+2q+1,\dots,\ i+3q-1\right\},$
@@
$C^{k}$ =$=\left\{i+kq,\ i+kq+1,\dots,i+(k+1)q-1\right\},$
@@
$C^{\alpha-1}$ =$=\left\{i+(\alpha-1)q,\ i+(\alpha-1)q+1,\dots,\ i-q\right\}.$
Note that the set $\left\{C^{1},\dots,C^{\alpha-1}\right\}$ forms a partition of $V(G\setminus i)$, i.e., $\bigcup_{k=1,\dots,\alpha-1}C^{k}=V(G\setminus i)$ and $C^{\ell}\cap C^{k}=\emptyset$, for all $\ell,k=1,\dots,\alpha-1$, with $\ell\not=k$. Hence, the sum of clique inequalities $\sum_{j\in C^{k}}x_{j}\leq 1$, for $k=1,\dots,\alpha-1$ yield the inequality
@@
Theorem 1 completes the theoretical picture for $N_{+}(\operatorname{QSTAB}(G))$ along the line drawn in [43]. In fact, all of the traditional combinatorial inequalities, namely, clique, odd cycle, odd antihole, odd wheel, web and antiweb inequalities turn out to be satisfied by $N_{+}(\operatorname{QSTAB}(G))$. It is interesting to mention that Giandomenico et al. [28] proved the same result for a non-compact linear relaxation obtained by the application of the $N(K,K)$ lifting operator by Lovász and Schrijver to $\operatorname{QSTAB}(G)$. The comparison between $N(K,K)$ and $N_{+}(\cdot)$ was indeed suggested by Lovász and Schrijver [43] as an interesting open issue.
Another insight of the progress of $N_{+}(\operatorname{QSTAB}(G))$ towards $N_{+}(\operatorname{FRAC}(G))$ stems from the analysis of graph $\hat{G}$ of Figure 1 which has been shown in [6] that $N_{+}(\operatorname{FRAC}(\hat{G}))$ gives rise to a non-polyhedral relaxation. Here we have $\lambda(\hat{G})=2.146$. Now, optimizing over $N_{+}(\operatorname{QSTAB}(\hat{G}))$ one gets $\mu(\hat{G})=2$. This implies that $N_{+}(\operatorname{QSTAB}(\hat{G}))$ is actually polyhedral.
@@
From a computational perspective, a major issue of $N_{+}(\operatorname{QSTAB}(G))$ is the linear description of $\operatorname{QSTAB}(G)$ that, in general, contains exponentially many inequalities. However, as ${\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$, we still have ${M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$ and one can consider the convex set:
@@
Note that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ still implies McCormick inequalities (14)–(17). Remarkably, it can be built in polynomial time as a suitable set $\cal C$ can be determined by a greedy algorithm (e.g., the one proposed in Letchford et al. [40]), and the number of inequalities describing $\operatorname{QSTAB}(G,\mathcal{C})$ is bounded by $|E|$ (for increasing graph density, it is also considerably smaller than $|E|$). Optimizing $\sum_{i\in V}x_{i}$ over $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ returns an upper bound that we denote as $\mu(G,{\cal C})$. We will experiment with such a relaxation in Section 6.

### v3

@@ 3.2.1 The Lovász theta relaxation
The Lovász theta relaxation [42] has been introduced in 1979, yet we refer to the presentation of Lovász and Schrijver [43]. Introduce the quadratic variable $X_{ij}$, representing the product $x_{i}x_{j}$ for all $\{i,j\}\in V\times V$ and let $X=xx^{\top}$ be the associated symmetric matrix of order $n=|V|$. An upper bound on $\alpha(G)$ is given by
$\theta(G)=\max$ $\;\sum_{i\in V}x_{i}$
s.t. $\;X_{ii}=x_{i},\quad i\in V,$ (SDP-$\theta$)
@@
$\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\succeq 0.$
The projection $\operatorname{TH}(G)$ of the feasible region of (SDP-$\theta$) onto the subspace of the $x_{i}$ variables is known as the theta body, a convex but not a polyhedral set in general. Remarkably, Grötschel et al. [33] proved that $\operatorname{STAB}(G)\subseteq\operatorname{TH}(G)\subseteq{\operatorname{QSTAB}(G)}$, where equality holds if and only if $G$ is perfect. This implies that there exists a polynomial-time separation algorithm for a class of inequalities which includes all clique inequalities. A thorough comparison between the original formulation in [42] and (SDP-$\theta$) is provided by Galli and Letchford [26]. $\theta(G)$ can be computed in polynomial time up to an arbitrary precision and often provides a strong bound to the stability number $\alpha(G)$, typically better than those obtained from linear relaxations see, e.g., [28, 30, 38, 69]. Another interesting feature of (SDP-$\theta$) concerns its computational behavior. In fact, (SDP-$\theta$) is handled better than similarly sized unstructured SDPs by general SDP solvers. Moreover, several algorithms have been designed to solve it efficiently as in, e.g., [44, 54, 29].
Therefore, $\theta(G)$ achieves a good compromise between strength of the upper bound and computational burden. From a theoretical viewpoint, Busygin and Pasechnik [13] showed that, unless P=NP, no polynomially computable upper bound to $\alpha(G)$ which is provably smaller than $\theta(G)$ can be found. However, several stronger relaxations have been obtained by adding linear inequalities to (SDP-$\theta$), as reviewed in the following subsections.
@@ 3.2.2 The Schrijver relaxation
Schrijver [59] observed that the PSD condition in (SDP-$\theta$) does not imply the non-negativity on $X$. Hence, by adding the inequalities
$X_{ij}\geq 0,\quad\{i,j\}\in\bar{E}$ (SDP-$\theta^{+}$a)
to (SDP-$\theta$), one yields a model that we denote as (3.2.2) and the corresponding upper bound as $\theta^{+}(G)$. Accordingly, $\operatorname{TH}^{+}(G)$ denotes the convex set obtained by its projection onto the subspace of $x_{i}$ variables.
@@ 3.2.3 The Lovász-Schrijver relaxation
The Lovász-Schrijver relaxation for the SSP is obtained by applying the operator $N_{+}(\cdot)$ to $\operatorname{FRAC}(G)$. $N_{+}(\operatorname{FRAC}(G))$ corresponds to (SDP-$\theta$) plus the following linear inequalities:
$X_{ij}$ $\geq 0,$ $\{i,j\}\in\bar{E},$ (11a)
$X_{ik}+X_{jk}$ $\leq x_{k},$ $\{i,j\}\in E,k\neq i,j,$ (11b)
$x_{i}+x_{j}+x_{k}$ $\leq 1+X_{ik}+X_{jk},$ $\{i,j\}\in E,k\neq i,j.$ (11c)
Lovász and Schrijver [43] assessed the theoretical strength of $N_{+}(\operatorname{FRAC}(G))$ by showing that it satisfies several well-known classes of valid inequalities for $\operatorname{STAB}(G)$. Namely, they proved that $N_{+}(\operatorname{FRAC}(G))$ satisfies all clique, odd cycle, odd antihole and odd wheel inequalities. Giandomenico and Letchford [27] showed that, in addition, it satisfies all web inequalities. Optimizing $\sum_{i\in V}x_{i}$ over $N_{+}(\operatorname{FRAC}(G))$ returns an upper bound on $\alpha(G)$ that we denote as $\lambda(G)$. From a more general perspective, it is interesting to observe the key role of the positive semidefiniteness constraint. In fact, Burer and Vandenbussche [12] compared $\lambda(G)$ with the optimal value of the corresponding LP relaxation obtained by the Sherali-Adams procedure, observing a substantial improvement due to it.
Handling such a strong relaxation in practice is not straightforward. The first related computational results, obtained by a lift-and-project cutting plane method, appear in [3]. Results on optimizing directly over $M_{+}(\operatorname{FRAC}(G))$ are illustrated in Dash [21] and, more recently, in Burer and Vandenbussche [12] using a specialized augmented Lagrangian method. The resulting upper bounds on $\alpha(G)$ may be considerably better than $\theta^{+}(G)$ but at the expense of a large increase in running times. In practice, this relaxation turns out to be hardly tractable for graphs of medium/large size.
@@ 3.2.4 Further SDP relaxations
Variants of the previous relaxations are presented in Dukanovic and Rendl [23] and in Gruber and Rendl [34]. In the former, extra triangle inequalities are added to $N_{+}(\operatorname{FRAC}(G))$ to obtain a very strong relaxation. In the latter, relaxations are specialized with respect to graph density to improve efficiency. We refer to Galli and Letchford [26] for an exhaustive review of these methods. Two SDP-related methodologies are also presented in the literature. In Gaar and Rendl [24], strong SDP relaxations of several combinatorial optimization problems are investigated by exploiting the exact subgraph inequalities. Despite the power of this technique, the experience reveals that strengthening the natural SDP relaxation of the SSP is typically harder than other problems, e.g., max-cut, and confirms that improving the $\theta$ bound towards $\alpha(G)$, even by a small amount, is actually challenging. Another effective method (presented in the context of the clique number) based on adding non-valid inequalities has been developed in Locatelli [41]. This turns out to achieve quite strong bounds for both structured and random graphs with up to 150 vertices.
@@ 4 New SDP relaxations
$x_{i}x_{j}\rightarrow$ $\quad X_{ij}\geq 0,$ (12a)
$x_{i}(1-x_{j})\rightarrow$ $\quad x_{i}\geq X_{ij},$ (12b)
$(1-x_{i})x_{j}\rightarrow$ $\quad x_{j}\geq X_{ij},$ (12c)
$(1-x_{i})(1-x_{j})\rightarrow$ $\quad x_{i}+x_{j}-X_{ij}\leq 1.$ (12d)
The linear inequalities (12a)–(12d) are known as McCormick [46] inequalities and describe the convex hull of the set $\{(x_{i},x_{j},y_{ij})\in\{0,1\}^{3}:y_{ij}=x_{i}x_{j}\}$. It is easy to see that they are not implied by the PSD condition (6) and should be explicitly added to the relaxations illustrated below unless implied by other constraints.
@@ 4.1 Relaxation $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$
In order to investigate $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ we first describe the application of the operator $M_{+}(\cdot)$ to $\operatorname{QSTAB}(G)$. The lifting step consists in multiplying each clique inequality $\sum_{i\in C}x_{i}\leq 1$, $C\in{\cal K}$, by $x_{j}$ and $(1-x_{j})$, for all $j\in V$. Then, adding the condition $X_{ii}=x_{i}$, one obtains
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C\setminus\{j\}}X_{ij}\leq 0,$ if $j\in C$, (13a)
$x_{j}\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum_{i\in C}X_{ij}-x_{j}\leq 0,$ if $j\not\in C$, (13b)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}x_{i}-\sum\limits_{i\in C\setminus\{j\}}X_{ij}\leq 1,$ if $j\in C$, (13c)
$(1-x_{j})\cdot\left(\sum_{i\in C}x_{i}\leq 1\right)\rightarrow$ $\sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}\leq 1,$ if $j\not\in C$. (13d)
Inequalities (13b), called clique-variable inequalities, have been introduced in [28]. One can observe that
@@
    Constraints (13a) imply $X_{ij}=0$ for all $\{i,j\}\in E$ and constraints (13c) reduce to the corresponding clique inequality $\sum_{i\in C}x_{i}\leq 1,C\in{\cal K}$, which is already implied from (13b) and (13d) by Remark 1;
@@
    $X_{ij}=0$ for all $\{i,j\}\in E$, $X_{ij}\geq 0$ for all $\{i,j\}\in\bar{E}$ and constraints (13b) imply inequalities (12b) and (12c);
@@
$\begin{split}{M_{+}(\operatorname{QSTAB}(G))}=\left\{\begin{pmatrix}1&x^{\top}\\ x&X\end{pmatrix}\succeq 0:\right.&\\ \sum\limits_{i\in C}X_{ij}-x_{j}&\leq 0,\quad C\in\mathcal{K},j\in V\setminus C,\\ \sum\limits_{i\in C}(x_{i}-X_{ij})+x_{j}&\leq 1,\quad C\in\mathcal{K},j\in V\setminus C,\\ x_{i}&=X_{ii},\quad i\in V,\\ X_{ij}&=0,\quad\{i,j\}\in E,\\ X_{ij}&\geq 0,\quad\{i,j\}\in{\bar{E}}\\ x_{i}&\geq 0,\quad i\in V\bigg\}.\end{split}$
Moreover, one has
@@
We denote the resulting upper bound on $\alpha(G)$ by
$\mu(G)=\max\left\{\sum_{i\in V}x_{i}:\begin{pmatrix}1&x^{\top}\cr x&X\cr\end{pmatrix}\in{M_{+}(\operatorname{QSTAB}(G))}\right\}$
The following example shows that the containment of Theorem 1 can be strict.
@@
The example indeed suggests a general result. In fact, we can show that, unlike $N_{+}(\operatorname{FRAC}(G))$, the relaxation $N_{+}(\operatorname{QSTAB}(G))$ implies the class of antiweb inequalities.
@@
$N_{+}(\operatorname{QSTAB}(G))$ implies antiweb inequalities.
Proof Let $G$ be the $(p,q)$-antiweb graph with $V=\left\{1,\ldots,p\right\}$. In what follows, we use arithmetic modulo $p$. Note that, by definition, any consecutive $q$ nodes in $G$ form a clique. By denoting $\alpha=\left\lfloor p/q\right\rfloor$, the $(p,q)$-antiweb inequality is
@@
Given any $i\in V$, we denote by $G\setminus i$ the subgraph induced by the vertex set
@@
is valid for $\operatorname{QSTAB}(G\setminus i)$}, for all $i\in V$, then the theorem’s statement will follow from Lemma 1. Let $i$ be any node in $V$ and consider the following $\alpha-1$ cliques in $G\setminus i$
$C^{1}$ $=\left\{i+q,\ i+q+1,\dots,\ i+2q-1\right\},$
$C^{2}$ $=\left\{i+2q,\ i+2q+1,\dots,\ i+3q-1\right\},$
@@
$C^{k}$ $=\left\{i+kq,\ i+kq+1,\dots,i+(k+1)q-1\right\},$
@@
$C^{\alpha-1}$ $=\left\{i+(\alpha-1)q,\ i+(\alpha-1)q+1,\dots,\ i-q\right\}.$
Note that the set $\left\{C^{1},\dots,C^{\alpha-1}\right\}$ forms a partition of $V(G\setminus i)$, i.e.,
$\bigcup_{k=1,\dots,\alpha-1}C^{k}=V(G\setminus i)$
and $C^{\ell}\cap C^{k}=\emptyset$, for all $\ell,k=1,\dots,\alpha-1$, with $\ell\not=k$. Hence, the sum of clique inequalities $\sum_{j\in C^{k}}x_{j}\leq 1$, for $k=1,\dots,\alpha-1$ yields the inequality
@@
Theorem 1 completes the theoretical picture for $N_{+}(\operatorname{QSTAB}(G))$ along the line drawn by Lovász and Schrijver [43]. In fact, all of the traditional combinatorial inequalities, namely, clique, odd cycle, odd antihole, odd wheel, web and antiweb inequalities turn out to be satisfied by $N_{+}(\operatorname{QSTAB}(G))$. It is interesting to mention that Giandomenico et al. [28] proved the same result for a non-compact linear relaxation obtained by the application of the $N(K,K)$ lifting operator by Lovász and Schrijver to $\operatorname{QSTAB}(G)$. The comparison between $N(K,K)$ and $N_{+}(\cdot)$ was indeed suggested by Lovász and Schrijver [43] as an interesting open issue.
Another insight of the progress of $N_{+}(\operatorname{QSTAB}(G))$ towards $N_{+}(\operatorname{FRAC}(G))$ stems from the analysis of graph $\hat{G}$ of Figure 1 which has been shown in [6] that $)N_{+}(\operatorname{FRAC}(\hat{G}$)) gives rise to a non-polyhedral relaxation. Here we have $\lambda(\hat{G})=2.146$. Now, optimizing over $)N_{+}(\operatorname{QSTAB}(\hat{G}$)) one gets the stability number $\mu(\hat{G})=\alpha(\hat{G})=2$. This may suggest that $)N_{+}(\operatorname{QSTAB}(\hat{G}$)) is actually polyhedral.
@@
From a computational perspective, a major issue of $N_{+}(\operatorname{QSTAB}(G))$ is that the linear description of $\operatorname{QSTAB}(G)$ , in general, contains exponentially many inequalities. To get rid of this issue, we consider $\operatorname{QSTAB}(G,\mathcal{C})$}, for any edge clique cover $\cal C\in\cal Q$. The application of the lifting operator to this relaxation yields the convex set
@@
Since ${\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$, we have ${M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))}\subseteq{M_{+}(\operatorname{FRAC}(G))}$. Also note that $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ still implies McCormick inequalities (12a)–(12d). Remarkably, it can be built in polynomial time as a minimal edge clique cover ${\cal C}\in{\cal Q}$ can be determined by a greedy algorithm (e.g., the one proposed in Letchford et al. [40]), and the number of inequalities describing $\operatorname{QSTAB}(G,\mathcal{C})$ is at most $|E|$ (for increasing graph density, it is also considerably smaller than $|E|$). Optimizing $\sum_{i\in V}x_{i}$ over $M_{+}(\operatorname{QSTAB}(G,\mathcal{C}))$ returns an upper bound that we denote as $\mu(G,{\cal C})$. We will experiment with such a relaxation in Section 6.